a certain 90-mile trip took 2 hours. exactally 1/3 of the distance traveledwas by raail, and this part of the trip took 1/5 of the travel time. what was the average rate, in miles per hour of the rail portion of the trip?
Answer by nerdybill(7384)

a certain 90-mile trip took 2 hours. exactally 1/3 of the distance traveledwas by raail, and this part of the trip took 1/5 of the travel time. what was the average rate, in miles per hour of the rail portion of the trip?
.
1/3 of the trip is then:
(1/3)(90) = 30 miles
.
Time it takes to travel that distance is:
(1/5)(2) = 2/5 of an hour
.
Average rate:
30/(2/5) = 150/2 = 75 mph